OpenStudy (hba):

if x+iy=(3+2i)(3-2i) then x and y are equal to

OpenStudy (anonymous):

expand RHS first

hartnn (hartnn):

well did u get y= 0 ?

OpenStudy (hba):

Wait I am telling you the answer @hartnn

OpenStudy (hba):

x+iy=9-4i^2

OpenStudy (anonymous):

then x=9

hartnn (hartnn):

umm....i^2=-1

hartnn (hartnn):

so x= ?

OpenStudy (hba):

yeah ^

OpenStudy (anonymous):

x=9 iy=-4i^(2) calculate it

hartnn (hartnn):

what did u get x as ?

OpenStudy (hba):

x is 9 and y is 4

hartnn (hartnn):

nopes

OpenStudy (hba):

Or x+iy=13

hartnn (hartnn):

9-4(-1) = 13 +0i =x+iy

OpenStudy (hba):

Yeah right Thanks @kryton1212 Confused me :)

OpenStudy (anonymous):

lol...

hartnn (hartnn):

u got it @kryton1212 ?

OpenStudy (anonymous):

got it-.-

OpenStudy (hba):

x=13 and y=0 :)

OpenStudy (hba):

Intelligent @kryton1212